# How Vulnerable is India’s Economy to Foreign Sanctions?

### 12 Mar 2026 v1

@@ Abstract
This paper develops a simple model of the world supply chain to estimate the effects of sanctions that restrict the flow of inputs from one country to another. Such restrictions operate through changes in the weights of the global production network: the sanctioning country ceases supplying certain inputs to the target country and reallocates its production to other destinations. Using the OECD Inter-Country Input–Output tables, we calibrate the model to assess the vulnerability of the Indian economy. We consider two classes of counterfactuals: restrictions on a single sector of a foreign country supplying India, and restrictions on all sectors of a foreign country supplying India. We then rank foreign countries and foreign country-sectors by the risk that their supply restrictions pose to economic activity in India. Our results show that India’s greatest country-level vulnerability is to Saudi Arabia, followed by the United Arab Emirates, China, Singapore, the United States, and Russia.
@@ 1 Introduction
One measure of interdependence between national economies is the ratio of world exports to world GDP. In 1830, this ratio was about $6\%$; today it is about $30\%$ (Federico and Tena-Junguito, 2019). Another measure of international production interdependence is the foreign value-added content of exports, or the share of exports that embodies imported intermediate inputs. By this metric too, cross-country production linkages increased markedly after 1990 (World Bank, 2026a, 2020). Yet such measures of dependence likely understate the vulnerability of national economies to disruptions in international input flows. In several strategically important inputs, global supply is heavily concentrated in a small number of countries: China is the dominant exporter of natural graphite, Chile accounts for roughly two-thirds of lithium carbonate exports, and the Democratic Republic of the Congo remains overwhelmingly important in cobalt (World Bank, 2026c, b; International Energy Agency, 2025). In such a skewed setting, we must be able to study the consequences of restrictions imposed by a small number of countries—or even a single country—on the export of specific inputs.
@@
Our empirical analysis focuses on the vulnerability of the Indian economy, although none of the theoretical apparatus developed in the paper is specific to her. In the country-sector counterfactuals, where a foreign country ceases exporting the inputs of one particular sector to India, the Indian economy is most vulnerable to the oil and gas sectors of Saudi Arabia and the United Arab Emirates, followed by Australia’s coal mining and China’s chemicals. Indeed, nearly all of the top twenty entries, out of more than 3,500 country-sectors, belong to fossil fuels, mining, and related upstream activities, with Singapore’s water transport sector being the primary exception. No European country-sector appears in the top twenty. This pattern is broadly preserved when we turn to country-wide sanctions, in which all sectors of a foreign country cease supplying intermediate inputs to India. Saudi Arabia, the United Arab Emirates, and China remain the three largest sources of vulnerability. The main difference is that advanced European economies now enter the ranking, including the United Kingdom, Germany, Denmark; Japan too makes it to the list.
@@ 4.3 Country-sector vulnerability
Figure 2 ranks the foreign country-sectors that generate the greatest vulnerability for the Indian economy. The two largest vulnerabilities arise from the extraction of petroleum and natural gas in Saudi Arabia and the United Arab Emirates, followed closely by coal mining in Australia and chemicals in China. Figure 3 reports the same ranking after excluding oil and gas production. The broad pattern remains unchanged: India’s economy appears especially vulnerable to foreign supplies of chemicals, ores, and other natural resources, rather than to high-end technology sectors. Korea and Japan appear in the ranking, but not as providers of frontier technology; instead, they matter through a more basic upstream sector, i.e. ‘iron and steel’. It is also striking that sectors from the major European economies are largely absent from the very top of the ranking.
Figure 4 displays the distribution of India’s vulnerability across foreign country-sectors. Note that the vulnerability index plotted on a logarithmic horizontal axis, which makes makes the extreme skewness of the distribution apparent: a very small number of foreign country-sectors account for a disproportionate share of India’s vulnerability, while the vast majority contribute very little. Indeed, the top 10 country-sectors account for about 33% of total vulnerability, and the top 100 for nearly 75%, out of a universe of more than 3,500 country-sectors.
@@
Figure 3: India’s vulnerability to the closure of intermediate-input flows from specific foreign country-sectors, excluding sectors that produce oil and gas.
@@ 4.4 Country vulnerability
Figure 5 presents India’s vulnerability at the country level, listing the twenty foreign countries whose complete interruption of intermediate-input flows would be most disruptive for the Indian economy. Unsurprisingly, Saudi Arabia and the United Arab Emirates appear at the top of the ranking, consistent with the country-sector results in Section 4.3, where India’s dependence on petroleum and natural gas from these economies emerged clearly. They are followed by China, which is likewise unsurprising given the wide range of intermediate goods it supplies to the world economy, including India. Unlike the country-sector ranking, however, the country-level ranking also brings advanced European economies such as Germany and Denmark into prominence.
Figure 6 complements this ranking with a world map in which countries are shaded by the vulnerability they pose to the Indian economy under a country-wide restriction, that is, under the complete cessation of intermediate-input flows from all of their sectors to India. The geographic pattern confirms that India’s external vulnerability is concentrated in a relatively small set of countries, with especially strong exposure to energy and ore suppliers. In fact, the top ten countries account for more than 70% of total vulnerability. More generally, India’s exposure is far from uniform across trading partners: some countries matter very little in aggregate, whereas others are capable of generating substantial disruption if they curtail the flow of intermediate inputs.
@@
Figure 6: Country-level vulnerability of the Indian economy under country-wide supply restrictions.

### 14 Mar 2026 v2

@@ Abstract
This paper develops a simple model of the world supply chain to estimate the effects of sanctions that restrict the flow of inputs from one country to another. Such restrictions operate through changes in the weights of the global production network: the sanctioning country ceases supplying certain inputs to the target country and reallocates its production to other destinations. Using the OECD Inter-Country Input–Output tables, we calibrate the model to assess the vulnerability of the Indian economy. We consider two classes of counterfactuals: restrictions on a single sector of a foreign country supplying India, and restrictions on all sectors of a foreign country supplying India. We then rank foreign countries and foreign country-sectors by the risk that their supply restrictions pose to economic activity in India. Our results show that India’s greatest country-level vulnerability is to China, followed by the United Arab Emirates, the United States, Saudi Arabi and Russia, with the vulnerability to China being twice as much that to the UAE.
@@ 1 Introduction
One measure of interdependence between national economies is the ratio of world exports to world GDP. In 1830, this ratio was about $6\%$; today it is about $30\%$. Another measure of international production interdependence is the foreign value-added content of exports, or the share of exports that embodies imported intermediate inputs. By this metric too, cross-country production linkages increased markedly after 1990. Yet such measures of dependence likely understate the vulnerability of national economies to disruptions in international input flows. In several strategically important inputs, global supply is heavily concentrated in a small number of countries: China is the dominant exporter of natural graphite, Chile accounts for roughly two-thirds of lithium carbonate exports, and the Democratic Republic of the Congo remains overwhelmingly important in cobalt [^1] For data on graphite, lithium and other critical minerals see World Bank (2026c, b); International Energy Agency (2025). For exports and trade related statistics, see (Federico and Tena-Junguito, 2019; World Bank, 2026a, 2020) . In such a skewed setting, we must be able to study the consequences of restrictions imposed by a small number of countries—or even a single country—on the export of specific inputs.
@@
Our empirical analysis focuses on the vulnerability of the Indian economy, although none of the theoretical apparatus developed in the paper is specific to her. In the country-sector counterfactuals, where a foreign country ceases exporting the inputs of one particular sector to India, the Indian economy is most vulnerable to the oil and gas sectors of Saudi Arabia and the United Arab Emirates, followed by Australia’s coal mining and China’s chemicals. Indeed, nearly all of the top twenty entries, out of more than 3,500 country-sectors, belong to fossil fuels, mining, and related upstream activities, with Singapore’s water transport sector being the primary exception. No European country-sector appears in the top twenty. This pattern is broadly preserved when we turn to country-wide sanctions, in which all sectors of a foreign country cease supplying intermediate inputs to India. Saudi Arabia, the United Arab Emirates, and Russia figure in the top five, with China leading the group by far and the United States ranking third. The main difference between country-sector vulnerabilities and country vulnerabilities is that now some advanced European economies enter the ranking, including the United Kingdom, Germany, and France; Japan too makes it to the list.
@@ 4.3 Country-sector vulnerability
Figure 2 ranks the foreign country-sectors that generate the greatest vulnerability for the Indian economy. The two largest vulnerabilities arise from the extraction of petroleum and natural gas in Saudi Arabia and the United Arab Emirates, followed closely by coal mining in Australia and chemicals in China. Figure 3 reports the same ranking after excluding oil & gas production and mining. The broad pattern remains unchanged: India’s economy appears especially vulnerable to foreign supplies of chemicals and metals, rather than to high-end technology sectors. There are exceptions: the most prominent being ‘professional, scientific, and technical activities’ of the United States. It is striking that sectors from the most European economies are largely absent, even after removing oil, gas, and mining.
Figure 4 displays the distribution of India’s vulnerability across foreign country-sectors. Note that the vulnerability index plotted on a logarithmic horizontal axis, which makes makes the extreme skewness of the distribution apparent: a very small number of foreign country-sectors account for a disproportionate share of India’s vulnerability, while the vast majority contribute very little. Indeed, the top 10 country-sectors account for about a third of total vulnerability, and the top 100 for nearly three-quarters. This in a universe of more than 3,500 country-sectors.
@@
Figure 3: India’s vulnerability to the closure of intermediate-input flows from specific foreign country-sectors, excluding sectors that produce oil, gas, and mining.
@@ 4.4 Country vulnerability
Figure 5 presents India’s vulnerability at the country level, listing the twenty foreign countries whose complete interruption of intermediate-input flows would be most disruptive for the Indian economy. Unsurprisingly, China leads the ranking, followed by the United Arab Emirates, the United States, Saudi Arabia, and Russia. This is somewhat consistent with the country-sector results in Section 4.3, where India’s dependence on petroleum and natural gas from these economies (except China) emerged clearly. Unlike the country-sector ranking, however, the country-level ranking also brings advanced European economies such as Germany, the United Kingdom and France into the fray. Japan almost makes it to the top ten.
Figure 6 complements this ranking with a world map in which countries are shaded by the vulnerability they pose to the Indian economy under a country-wide restriction, that is, under the complete cessation of intermediate-input flows from all of their sectors to India. The geographic pattern confirms that India’s external vulnerability is concentrated in a relatively small set of countries, with especially strong exposure to energy and ore suppliers (with China being the exception). In fact, the top ten countries account for nearly 70% of total vulnerability. More generally, India’s exposure is far from uniform across trading partners: some countries matter very little in aggregate, whereas others are capable of generating substantial disruption if they curtail the flow of intermediate inputs.
@@
Figure 6: Vulnerability of the Indian economy under country-wide supply restrictions.
@@ 5 Concluding Thoughts
We have therefore presented a rudimentary analysis of a problem that cannot be wished away, at least not for as long as human society relies on the organizing capacity of political forces. While the mathematical framework we develop is quite general, the empirical analysis is only as good as the data on which it rests, which in our case are the OECD input-output tables. It is no simple matter to construct reliable sectoral flow data across countries. Many countries do not regularly publish high-quality information on the sectoral composition of output, still less on the pattern of flows across sectors and borders. In such cases, the OECD must rely, at least in part, on structural assumptions and indirect methods that use what is known about some economies to infer what is less well measured in others. This is not a defect peculiar to our exercise, but a limitation that attaches more generally to any attempt to study the global production network. Our empirical results should therefore be read in that spirit: not as precise point estimates but as rough approximations.